2) if a+bi=2+i / 1+i what is the value of a+b
A) a=1 b=3
B) a=-1/2 b=-3/2
C)a=3/2 b=-1/2
D)a=1/2 b=-3/2

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

(2+i)/(1+i); multiply denominator by conjugate (1-i) and that becomes 1-i^2=2
numerator is (2+i)(1-i)=2-i-i^2=3-i
fraction is (3-i)/2= (3/2)-(1/2)i
C.

Answer by ikleyn(53763)

.
On complex numbers, see the relevant lessons
    - Complex numbers and arithmetical operations on them
    - Complex plane
    - Addition and subtraction of complex numbers in complex plane
    - Multiplication and division of complex numbers in complex plane
    - Raising a complex number to an integer power

    - Solved problems on taking roots of complex numbers
    - Solved problems on arithmetic operations on complex numbers
in this site.

Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic  "Complex numbers".